OpenStudy (anonymous):

graph the function. compare with graph of y=x^2 ------>1. y=-x^2+2 graph includedhelp

OpenStudy (sburchette):

If the coefficient of the x^2 term is greater than 1, it will cause the parabola to be more narrow. If the coefficient is less than 1, (but not equal to zero of course) it will cause the parabola to widen. This is also true for negative coefficients, but the parabola will also be flipped. So y=3x^2 will be more narrow than y=x^2. In the attachment, the black curve is y=x^2 and the red curve is y=3x^2.

OpenStudy (anonymous):